SuicidalSheep32:

Find the circumference and area of a circle with a diameter of 10 inches. Leave your answers in terms of pi.

Hero:

@SuicidalSheep32 what are the formulas for area and circumference of a circle?

SuicidalSheep32:

The area of a circle is Pi*r^2 Circumference is 2*Pi*r

Hero:

Okay, great. Are you also able to recall the formula for determining the radius of a circle when given its diameter? @SuicidalSheep32

SuicidalSheep32:

The radius is just half of the diameter.

Hero:

Very good, so in this case, what would be the radius of the circle?

SuicidalSheep32:

5

Hero:

Bingo. So at this point, it seems to me that you have all the tools you need to find the circumference and area of the circle. Have you already computed these values? If so, please share.

SuicidalSheep32:

the area is 78.54 circumference is 31.42

Hero:

But it says to "Leave your answers in terms of pi".

SuicidalSheep32:

so Pi 10 and how would I do the other one 20 pi?

SuicidalSheep32:

area first then circumference

Hero:

Whatever the answers are, always remember to include units. When you post your answers you should do as you did before: The area of the circle is ..... ; The circumference of the circle is ...

Hero:

And remember to use the correct formulas for each computation.

SuicidalSheep32:

I forgot to put the answers but none of them ate what I got Find the circumference and area of a circle with a diameter of 10 inches. Leave your answers in terms of pi. C = 5π; A = 20π C = 10π; A = 20π C = 10π; A = 25π C = 5π; A = 25π

SuicidalSheep32:

are*

Hero:

You already posted the formulas above, but I'll post them again: The area, A, of a circle with radius , r, can be found by \(A = \pi r^2\) The circumference of a circle, C, can be found by \(C = 2 \pi r\) We know that the radius is 5 inches, which means \(r = 5\) so all we have to do is plug that value in to each formula: \[A = \pi (5)^2\]\[C = 2 \pi (5)\] If we simplify each of the expressions for \(A\) and \(C\) what values do we end up with?

SuicidalSheep32:

A=π 10 and for C i am not sure I think it is 20

Hero:

Can you explain why you think \(5^2\) means the same thing as \(5 + 5\) ?

Hero:

Then explain how you got \(20\) after multiplying \(2 \times 5\)

SuicidalSheep32:

wow like 4th-grade stuff and i am screwing it up it is 25

Hero:

Yes correct. Now what is the area of the circle @SuicidalSheep32 ?

SuicidalSheep32:

78.5

Hero:

Once again "Leave your answers in terms of pi"

SuicidalSheep32:

A= 25 π

Hero:

Finally! But remember to include the UNITS

SuicidalSheep32:

What units? what do you mean?

Hero:

Area is always expressed in square units. Units being the unit of measure which in this case is inches.

SuicidalSheep32:

Thank you @hero you are the best!

Hero:

But you haven't even confirmed your answers @SuicidalSheep32

SuicidalSheep32:

It is C = 10π INCHES; A = 25π INCHES

Hero:

Correct.

SuicidalSheep32:

Ok thank you hero